OpenStudy (anonymous):

What is the vertex form of the equation? y = x^2 – 6x + 8

OpenStudy (anonymous):

take middle coefficient, divide it by 2, square the result, then both add and subtract that after the middle term so (-6/2)^2=9 y=x^2-6x+9-9+8 factor the first part, it'll always be a perfect square y=(x-3)^2-9+8 simplify y=(x-3)^2-1